### v1

# Small-dimensional normed barrelled Spaces

Abstract.
We prove that there is a normed barrelled space with dimension $\mathrm{non}(\mathcal{M})$, which denotes the smallest cardinality of a non-meager subset of $\mathbb{R}$. Consequently, it is consistent with $\mathsf{ZFC}$ that there is a normed barrelled space with dimension $<\!\mathfrak{c}$. This answers a question of Sánchez Ruiz and Saxon.
This is a special case of a more general theorem: for every infinite cardinal $\kappa$, every Banach space with density character $\kappa$ contains a barrelled subspace with dimension $\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})$.
We also prove that if the dual of a Banach space does not contain $c_{0}$ or $\ell^{p}$ for any $p\geq 1$, then that spaces does not have a barrelled subspaces with dimension $<\!\mathrm{cov}(\mathcal{N})$, which denotes the smallest cardinality of a collection of Lebesgue null sets covering $\mathbb{R}$. In particular, it is consistent with $\mathsf{ZFC}$ that no classical Banach spaces contain barrelled subspaces with dimension $\mathfrak{b}$. This partly answers another question of Sánchez Ruiz and Saxon.
Key words and phrases:
barrelled spaces, Banach spaces, meager sets, null sets
2020 Mathematics Subject Classification:
46A08, 46B26, 03E17
The first author is supported in part by NSF grant DMS-2154229.
1. Introduction
A classical theorem of Mazur states that every infinite-dimensional Fréchet space has dimension at least $\mathfrak{c}=|\mathbb{R}|$ (see [2, p. 76] or [12]). In other words,
[…]
Here, and throughout the paper, the “dimension” of a topological vector space space always means its algebraic dimension.
Building in this direction, Sánchez Ruiz and Saxon authored a series of three papers, [15, 16, 17], in which they sought to characterize the smallest cardinality of other kinds of infinite-dimensional topological vector spaces. For example, they prove in [15] that there is a metrizable barrelled space of dimension $\mathfrak{b}$ (their example is a modification of Tweddle’s space from [23]), and that there is no infinite-dimensional metrizable barrelled space of dimension $<\!\mathfrak{b}$. In other words,
$\mathfrak{b}\,=\,\min\left\{\kappa\colon\text{there is a metrizable barrelled space of infinite dimension }\kappa\right\}.$
Here $\mathfrak{b}$ denotes the unbounding number, which is one of the classic cardinal characteristics of the continuum (see [1, 5]). These are cardinal numbers that are defined in some combinatorial or topological way, and that lie between $\aleph_{1}$ and $\mathfrak{c}$.
Sánchez Ruiz and Saxon define another such cardinal:
$\mathfrak{y}\,=\,\min\left\{\kappa\colon\text{there is a normed barrelled space of infinite dimension }\kappa\right\}.$
Their characterization of $\mathfrak{b}$ described above shows $\mathfrak{b}\leq\mathfrak{y}$, and any separable Banach space witnesses $\mathfrak{y}\leq\mathfrak{c}$. They ask in [15] and [16] whether $\mathfrak{y}$ might not simply be an alternative characterization of an already-known cardinal characteristic, much like the characterizations of $\mathfrak{b}$ and $\mathfrak{c}$ described above. The natural candidates are $\mathfrak{y}=\mathfrak{b}$, like the analogous cardinal for metrizable barrelled spaces, or $\mathfrak{y}=\mathfrak{c}$, like with Mazur’s result for Fréchet spaces. In this paper we show that $\mathfrak{y}$ is not provably equal to $\mathfrak{c}$ (i.e., it is consistent with $\mathsf{ZFC}$ that $\mathfrak{y}<\mathfrak{c}$), and we rule out some of the most natural techniques for showing $\mathfrak{y}=\mathfrak{b}$. The main result of Section 2 is:
  • $\circ$
    There is a barrelled subspace of $\ell^{1}$ with dimension $\mathrm{non}(\mathcal{M})$. (Theorem 2.3)
The cardinal number $\mathrm{non}(\mathcal{M})$ is defined as the smallest size of a meager subset of $\mathbb{R}$. Clearly $\aleph_{1}\leq\mathrm{non}(\mathcal{M})$ (i.e., non-meager sets are uncountable), and the Baire Category Theorem implies $\mathrm{non}(\mathcal{M})\leq\mathfrak{c}$. The cardinal $\mathrm{non}(\mathcal{M})$, known as the uniformity number of the meager ideal, is one of the classical cardinal characteristics of the continuum. Combining this with the results of [16],
$\aleph_{1}\,\leq\,\mathfrak{b}\,\leq\,\mathfrak{y}\,\leq\,\mathrm{non}(\mathcal{M})\,\leq\,\mathfrak{c}.$
It is consistent with $\mathsf{ZFC}$ that $\aleph_{1}=\mathfrak{c}$. This assertion, known as the Continuum Hypothesis and denoted $\mathsf{CH}$, implies all the above cardinals are equal. However, it is also consistent that $\mathrm{non}(\mathcal{M})<\mathfrak{c}$. (In fact this is true in Cohen’s original model showing that the failure of $\mathsf{CH}$ is consistent with $\mathsf{ZFC}$.) Thus, as a corollary to the theorem stated above, we obtain:
  • $\circ$
    It is consistent with $\mathsf{ZFC}$ that $\mathfrak{y}<\mathfrak{c}$. (Theorem 2.3, part 2)
The proof of Theorem 2.3 obtains a normed barrelled space with dimension $\mathrm{non}(\mathcal{M})$ not via some abstract construction, but by identifying a subspace of $\ell^{1}$ with size $\mathrm{non}(\mathcal{M})$ and with barrelled span. This suggests a natural generalization of the theorem, which turns out to be true: all Banach spaces have barrelled subspaces with “small” dimension. Proving this generalization is the topic of Section 3. The main results of Section 3 are:
  • $\circ$
    If $X$ is an infinite-dimensional Banach space with density character $\kappa$, there is a barrelled subspace of $X$ with dimension $\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})$. (Theorem 3.4)
  • $\circ$
    It is consistent with $\mathsf{ZFC}$ that $\mathfrak{c}$ is arbitrarily large, and that every Banach space with density character $<\!\mathfrak{c}$ has an infinite-dimensional barrelled subspace with dimension $<\!\mathfrak{c}$. (Corollary 3.6)
What exactly $\mathrm{cf}[\kappa]^{\omega}$ means is explained in Section 3, but let us point out for now that $\mathrm{cf}[\aleph_{0}]^{\omega}=1$, which means $\mathrm{cf}[\aleph_{0}]^{\omega}\cdot\mathrm{non}(\mathcal{M})=\mathrm{non}(\mathcal{M})$. Thus a special case of the above theorem is:
  • $\circ$
    Every infinite-dimensional separable Banach space has a barrelled subspace with dimension $\mathrm{non}(\mathcal{M})$.
The arguments in Section 3 follow the same pattern as those in Section 2, but with some additional technicalities not present in the case $X=\ell^{1}$. We have chosen to present both proofs because $(1)$ readers only interested in the special case can skip Section 3 altogether, and $(2)$ even those readers who do not with to skip Section 3 may find the arguments there easier to digest after first reading Section 2.
After proving these theorems about the existence of small-dimensional barrelled subspaces, in Section 4 we turn to the related question of how small is too small. Our main results in this section are:
  • $\circ$
    If the dual of a Banach space $X$ does not contain $c_{0}$ or $\ell^{p}$ for any $p\geq 1$, then $X$ does not have a barrelled subspace with dimension $<\!\mathrm{cov}(\mathcal{N})$. It is consistent they have no barrelled subspaces of dimension $\mathfrak{b}$. (Corollary 4.10)
This result is proved by finding a way to associate members of a Lebesgue probability space to norm-unbounded sequences of functionals on $X$, in such a way that any particular $\mathbf{x}\in X$ will almost surely not witness the norm-unboundedness of a “random” norm-unbounded sequence. For the $\ell^{1}$ case, these random norm-unbounded sequences are obtained by taking random walks in the dual of $X$.
The cardinal $\mathrm{cov}(\mathcal{N})$ is another cardinal characteristic of the continuum, known as the covering number of the null ideal and defined as
$\mathrm{cov}(\mathcal{N})\,=\,\textstyle\min\left\{|\mathcal{F}|\colon\mathcal{F}\text{ is a family of Lebesgue null subsets of }\mathbb{R}\text{ and }\bigcup\mathcal{F}\supseteq\mathbb{R}\right\}.$
Amongst the classical cardinal characteristics mentioned so far, the following inequalities are provable in $\mathsf{ZFC}$:
[…]
No further inequalities are provable: it is consistent with $\mathsf{ZFC}$ that all these inequalities are strict (even simultaneously), and either of the orderings $\mathfrak{b}<\mathrm{cov}(\mathcal{N})$ and $\mathrm{cov}(\mathcal{N})<\mathfrak{b}$ is consistent (though, of course, not simultaneously).
In particular, because the inequality $\mathfrak{b}<\mathrm{cov}(\mathcal{N})$ is consistent, our results show that it is consistent every space whose dual does not contain $c_{0}$ or any $\ell^{p}$ does not contain barrelled subspaces of dimension $\mathfrak{b}$.
If this were true for all Banach spaces, it would mean $\mathfrak{b}<\mathfrak{y}$. At present we do not know how to prove this for all Banach spaces, only those whose duals do not contain $c_{0}$ or any $\ell^{p}$. But let us point out that this is a very broad class of Banach spaces. It was conjectured from the early days of Banach space theory (see [11]) that every Banach space contains either $c_{0}$ or $\ell^{p}$ for some $p\geq 1$. This conjecture stood for decades, until finally being refuted by Tsirelson in [22]. (For a more modern and more extreme example, see [7].) At any rate, every “classical” Banach space, i.e., all the spaces of continuous functions, or differentiable functions or of integrable functions, all the Banach spaces used in functional analysis for the first forty years of Banach space theory, contains $c_{0}$ or $\ell^{p}$ for some $p\geq 1$. There is also a technical sense in which every “simply defined” Banach space contains $c_{0}$ or $\ell^{p}$ for some $p\geq 1$ (see [10]).
Because of this, our results show that one cannot prove $\mathfrak{b}=\mathfrak{y}$ without using more exotic, non-classical Banach spaces. It also means that one cannot prove that every infinite-dimensional Banach space has a barrelled subspace with dimension $\mathfrak{b}$. In other words, the results from Section 2 cannot be improved by replacing $\mathrm{cov}(\mathcal{M})$ with $\mathfrak{b}$. Thus, while we have not proved the consistency of $\mathfrak{b}<\mathfrak{y}$, we have ruled out some natural strategies for trying to show $\mathfrak{b}=\mathfrak{y}$.
In what follows, a “Banach space” may be taken as a (complete normed) vector space over either $\mathbb{R}$ or $\mathbb{C}$. We assume the reader is familiar with the basic theory of such spaces, and recommend [21] as a reference. We use some topology and descriptive set theory in what follows, our basic references for which are [6] and [9], but we assume less familiarity with these topics. Although we do prove some consistency results, no knowledge of forcing is needed, with just one exception, the proof of Corollary 3.6, and this can be skipped without losing the main thread of the paper.
2. A barrelled subspace of $\ell^{1}$ with dimension $\mathrm{non}(\mathcal{M})$
We adopt the convention that a sequence in a set $A$ is a function $\omega\to A$, where $\omega$ denotes the first countable ordinal $\omega=\{0,1,2,\dots\}=\mathbb{N}\cup\{0\}$. Adopting the usual notation for ordinals, we take $n=\{0,1,2,\dots,n-1\}$ for all $n\in\omega$. This is merely a convenience, which enables us to write $f\!\restriction\!n$ instead of $f\!\restriction\!\{0,1,2,\dots,n-1\}$. If $s$ is a function with $\mathrm{dom}(s)=n=\{0,1,2,\dots,n-1\}$, then we write $|s|=n$.
The space $\omega^{\omega}$ of all functions $\omega\to\omega$ is called the Baire space. The standard basis for the Baire space consists of sets of the form
$\langle\langle s\rangle\rangle\,=\,\left\{x\in\omega^{\omega}\colon x\!\restriction\!|s|=s\right\},$
where $s$ is a function $n\to\omega$. Observe that $\langle\langle t\rangle\rangle\subseteq\langle\langle s\rangle\rangle$ if and only if $t$ is an extension of $s$. The Baire space is separable and completely metrizable, making it an example of a Polish space. It is homeomorphic to the space $\mathbb{R}\setminus\mathbb{Q}$ of irrational numbers.
Let $\mathcal{I}$ denote the set of injective functions $\omega\to\omega$. This is a subspace of $\omega^{\omega}$, and in fact it is a $G_{\delta}$ subspace. To see this, observe that
[…]
is an open set for every $n\in\omega$, and this means that $\mathcal{I}=\bigcap_{n\in\omega}U_{n}$ is $G_{\delta}$. Recalling that a subspace of a completely metrizable space is itself completely metrizable if and only if it is $G_{\delta}$ [6, Theorems 4.3.23-24], this shows $\mathcal{I}$ is completely metrizable, hence Polish. In fact, by the Alexandrov-Urysohn characterization of the Baire space, $\mathcal{I}$ is homeomorphic to $\omega^{\omega}$ [9, Exercise 7.2G]. In context, we write just $\langle\langle s\rangle\rangle$, rather than $\langle\langle s\rangle\rangle\cap\mathcal{I}$, to denote the basic open neighborhood of $\mathcal{I}$ described by an injective function $s:n\to\omega$.
Let $\ell^{1}$ denote the space of absolutely summable sequences in $\mathbb{R}$, considered as a Banach space with the usual norm, denoted $||\mathbf{x}||\,=\sum_{n\in\omega}|\mathbf{x}(n)|$ for all $\mathbf{x}\in\ell^{1}$. If $f:\omega\to\omega$ is increasing, define $\mathbf{x}_{f}\in\ell^{1}$ by setting
$\mathbf{x}_{f}(i)\,=\,\begin{cases}\frac{1}{3^{n}}&\text{ if }i=f(n),\\ 0&\text{ if }i\notin\mathrm{image}(f).\end{cases}$
If $f$ is increasing then $\mathbf{x}_{f}$ is what is known as a dilation of the sequence $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$. If $f$ is bijective then $\mathbf{x}_{f}$ is instead a rearrangement of the sequence $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$. In any case, if $f\in\mathcal{I}$ then $\mathbf{x}_{f}\in\ell^{1}$ and $||\mathbf{x}_{f}||=\sum_{n\in\omega}\frac{1}{3^{n}}=\nicefrac{{3}}{{2}}$. Thus the map $f\mapsto\mathbf{x}_{f}$ naturally identifies $\mathcal{I}$ with a subset of $\ell^{1}$.
Recall that $\ell^{\infty}$ denotes the space of bounded sequences in $\mathbb{R}$, with the norm $||\mathbf{x}||=\sup_{n\in\omega}|\mathbf{x}(n)|$. This is the dual of $\ell^{1}$, so that every $\mathbf{x}\in\ell^{\infty}$ represents a linear functional $\ell^{1}\to\mathbb{R}$ defined by setting $\mathbf{x}(\mathbf{y})\,=\,\sum_{n\in\omega}\mathbf{x}(n)\mathbf{y}(n)$.
Barrelled spaces are locally convex topological vector spaces for which every barrel (absolutely convex, closed, absorbing set) is a neighborhood of zero. These spaces are important because such fundamental theorems as the Uniform Boundedness Principle and the Closed Graph and Open Mapping Theorems hold in them. See [21] for more information on barrelled spaces. In what follows, we use the following fact about barrelled spaces; the reader otherwise unfamiliar with barrelled spaces may take this as the definition of barrelled for the purposes of this paper:
  • $\circ$
    Let $X$ be a Banach space and $X^{\prime}$ its dual space. A subspace $Z$ of $X$ is barrelled if and only if for any norm-unbounded sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ in $X^{\prime}$, i.e. a sequence with $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$, there is some $\mathbf{x}\in Z$ such that $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{x})|=\infty$. (See, e.g., [21, Theorem 5.21 and Corollary 24.5].)
For each $i\in\omega$, let $\mathbf{e}_{i}$ denote the member of $\ell^{1}$ with $\mathbf{e}_{i}(n)=0$ whenever $n\neq i$ and $\mathbf{e}_{i}(i)=1$.
Lemma 2.1.
Suppose $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is a sequence in $\ell^{\infty}$ with $\sup_{m\in\omega}||\mathbf{y}_{m}||_{\infty}=\infty$, but $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{e}_{i})\big{|}<\infty$ for all $i\in\omega$. Then
$\left\{f\in\mathcal{I}\colon\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}=\infty\right\}$
is a co-meager subset of $\mathcal{I}$.
Proof.
For each $i\in\omega$, define $B_{i}=\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{e}_{i})\big{|}$, which is finite by hypothesis. Given an injective function $s:n\to\omega$ and given some $k\leq n$, define
$M_{k}^{s}\,=\,\textstyle\max\Big{\{}k,\frac{3}{2}B_{s(0)},\frac{3}{2}B_{s(1)},\dots,\frac{3}{2}B_{s(k-1)}\Big{\}}.$
Observe that the value of $M^{s}_{k}$ does not depend on $s(i)$ for any $i\geq k$. In particular, if $f:\omega\to\omega$ is injective and $k\leq m<n$, then $M^{f\restriction m}_{k}=M^{f\restriction n}_{k}$. In light of this, let $M^{f}_{k}$ denote the common value of $M^{f\restriction n}_{k}$ for all $n>k$.
Let us say that $f:\omega\to\kappa$ satisfies property ()(*)( ∗ ) at some $k\in\omega$ if
$(*):\quad\text{There is some $m\in\omega$ such that $\big{|}\mathbf{y}_{m}(\mathbf{e}_{f(k)})\big{|}\,>\,||\mathbf{y}_{m}||-1\,>\,k3^{k}M^{f}_{k}$.}$
In this case, we say that $m\in\omega$ witnesses property ()(*)( ∗ ) at $k$ for $f$. Roughly, ()(*)( ∗ ) says that there is some $m\in\omega$ such that $\big{|}\mathbf{y}_{m}(\mathbf{e}_{f(k)})\big{|}$ is very large compared to $M^{f}_{k}$, and is also close to the norm of $\mathbf{y}_{m}$.
For each $n\in\omega$, let
$U_{n}$ =$\,=\,\left\{f\in\mathcal{I}\colon f\text{ satisfies }(*)\text{ at$k$ for at least $n$ different values of $k$}\right\}$
=$\,=\,\textstyle\bigcup\left\{\langle\langle s\rangle\rangle\colon s\text{ satisfies }(*)\text{ at $k$ for at least $n$ different values of $k$}\right\}.$
This set is clearly open, and we claim it is also dense in $\mathcal{I}$.
Fix $n\in\omega$. To see that $U_{n}$ is dense in $\mathcal{I}$, fix an injective function $s:k\to\omega$ for some $k\in\omega$, so that $\langle\langle s\rangle\rangle$ is a basic open subset $\mathcal{I}$. Using our hypothesis that $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$, first pick some $m\in\omega$ such that $||\mathbf{y}_{m}||>k3^{k}M_{k}^{s}+1$. Then pick some $j\in\omega$ such that $|\mathbf{y}_{m}(\mathbf{e}_{j})|>||\mathbf{y}_{m}||-1$. Define $s^{\prime}:(k+1)\to\kappa$ by setting
$s^{\prime}(i)\,=\,\begin{cases}s(i)&\text{ if }i<k\\ j&\text{ if }i=k.\end{cases}$
Observe that $|\mathbf{y}_{m}(\mathbf{e}_{j})|>k3^{k}M^{s}_{k}>B_{s(i)}\geq|\mathbf{y}_{m}(\mathbf{e}_{s(i)})|$ for all $i<k$. It follows that $j\neq s(i)$ for any $i<k$, and therefore $s^{\prime}$ is injective. Furthermore, our choices of $m$ and $j$ ensure that $s^{\prime}$ satisfies property ()(*)( ∗ ) at $k=|s|$.
Now fix an injective function $s_{0}:k\to\omega$ for some $k\in\omega$. Applying the argument from the previous paragraph $n$ times in a row, we can obtain sequences $s_{1},s_{2},\dots,s_{n}$ such that for each $i<n$, $s_{i+1}\!\restriction\!|s_{i}|=s_{i}$ and $s_{i+1}$ satisfies ()(*)( ∗ ) at $|s_{i}|$. But then $s_{n}$ satisfies ()(*)( ∗ ) for at least $n$ different values of $k$, namely for $k=|s_{0}|,|s_{1}|,\dots,|s_{n-1}|$. Thus $\langle\langle s_{n}\rangle\rangle\subseteq U_{n}$. Also, because $s_{i+1}\!\restriction\!|s_{i}|=s_{i}$ for all $i<n$,
$\langle\langle s_{0}\rangle\rangle\supseteq\langle\langle s_{1}\rangle\rangle\supseteq\langle\langle s_{2}\rangle\rangle\supseteq\dots\supseteq\langle\langle s_{n}\rangle\rangle.$
Therefore $\langle\langle s_{n}\rangle\rangle\subseteq\langle\langle s_{0}\rangle\rangle\cap U_{n}$. As $\langle\langle s_{0}\rangle\rangle$ was an arbitrary basic open subset of $\mathcal{I}$, this shows $U_{n}$ is dense in $\mathcal{I}$.
Because each of the $U_{n}$ is an open dense subset of $\mathcal{I}$,
$X\,=\,\textstyle\bigcap_{n\in\omega}U_{n}\,=\,\left\{f\in\mathcal{I}\colon f\text{ satisfies }(*)\text{ at $k$ for infinitely many values of $k$}\right\}$
is a dense $G_{\delta}$ subset of $\mathcal{I}$. To finish the proof of the lemma, it suffices to show that if $f\in X$ then $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}=\infty$.
Fix $f\in X$, fix some $k\in\omega$ such that $f$ satisfies ()(*)( ∗ ) at $k$, and fix some $m\in\omega$ such that $m$ witnesses property ()(*)( ∗ ) at $k$ for $f$. Because $\mathbf{y}_{m}$ is linear,
$M^{f}_{k}$ $\,=\,\textstyle\max\Big{\{}k,\frac{3}{2}B_{f(0)},\frac{3}{2}B_{f(1)},\dots,\frac{3}{2}B_{f(k-1)}\Big{\}}$
$\,\geq\,\textstyle\sup\left\{\,\big{|}\mathbf{y}_{m}\big{(}\sum_{i<k}x_{i}\mathbf{e}_{f(i)}\big{)}\big{|}\colon\sum_{i<k}|x_{i}|\leq\frac{3}{2}\right\}$
$\,\geq\,\textstyle\big{|}\mathbf{y}_{m}\big{(}\sum_{i<k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}\big{|}.$
Using this fact, together with the triangle inequality, the inequalities in ()(*)( ∗ ), and the fact that $M^{f}_{k}\geq k$, we get the following lower bound on $\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}$:
$\big{|}\mathbf{y}_{m}(\mathbf{d}_{f})\big{|}$ $\,=\,\textstyle\big{|}\,\mathbf{y}_{m}\big{(}\sum_{i\in\omega}\mathbf{d}_{f}(f(i))\mathbf{e}_{f(i)}\big{)}\,\big{|}$
$\,=\,\textstyle\big{|}\,\mathbf{y}_{m}\big{(}\sum_{i\in\omega}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}\,\big{|}$
$\,=\,\textstyle\big{|}\,\mathbf{y}_{m}\big{(}\sum_{i<k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}+\mathbf{y}_{m}\big{(}\frac{1}{3^{k}}\mathbf{e}_{f(k)}\big{)}+\mathbf{y}_{m}\big{(}\sum_{i>k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}\,\big{|}$
$\,\geq\,\textstyle\big{|}\mathbf{y}_{m}(\frac{1}{3^{k}}\mathbf{e}_{f(k)})\big{|}-\big{|}\mathbf{y}_{m}\big{(}\sum_{i<k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}\big{|}-\big{|}\mathbf{y}_{m}\big{(}\sum_{i>k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{)}\big{|}$
$\,\geq\,\textstyle\frac{1}{3^{k}}\big{|}\mathbf{y}_{m}(\mathbf{e}_{f(k)})\big{|}-M^{f}_{k}-||\mathbf{y}_{m}||_{\infty}\big{|}\big{|}\sum_{i>k}\frac{1}{3^{i}}\mathbf{e}_{f(i)}\big{|}\big{|}_{1}$
$\,\geq\,\textstyle\frac{1}{3^{k}}\big{|}\mathbf{y}_{m}(\mathbf{e}_{f(k)})\big{|}-M^{f}_{k}-\frac{1}{2\cdot 3^{k}}||\mathbf{y}_{m}||_{\infty}$
$\,\geq\,\textstyle\frac{1}{3^{k}}\Big{(}3^{k}\big{|}\mathbf{y}_{m}(\mathbf{e}_{f(k)})\big{|}-3^{k}M^{f}_{k}-\frac{1}{2}||\mathbf{y}_{m}||_{\infty}\Big{)}$
$\,\geq\,\textstyle\frac{1}{3^{k}}\Big{(}3^{k}(||\mathbf{y}_{m}||_{\infty}-1)-\frac{1}{k}(||\mathbf{y}_{m}||_{\infty}-1)-\frac{1}{2}||\mathbf{y}_{m}||_{\infty}\Big{)}$
$\,\geq\,\textstyle\frac{1}{3^{k}}\Big{(}(3^{k}-1)||\mathbf{y}_{m}||_{\infty}-3^{k}\Big{)}$
$\,\geq\,\textstyle\big{(}1-\frac{1}{3^{k}}\big{)}M^{f}_{k}-1$
$\,\geq\,\textstyle\frac{2}{3}k-1.$
Because $f\in X$, this shows that there are arbitrarily large values of $k$ such that $\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}\geq\frac{2}{3}k-1$ for some $m\in\omega$. Hence $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}=\infty$. ∎
Theorem 2.2.
If $X$ is a non-meager subset of $\mathcal{I}$, then
$Z\,=\,\mathrm{Span}\big{(}\!\left\{\mathbf{e}_{i}\colon i\in\omega\right\}\cup\big{\{}\mathbf{x}_{f}:\,f\in X\big{\}}\big{)}$
is a barrelled subspace of $\ell^{1}$.
Proof.
To show $Z$ is barrelled, it suffices to show that for any sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ in $\ell^{\infty}$, if $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is pointwise bounded on $Z$ then it is bounded in norm. We show the contrapositive: if a sequence is not bounded in norm, it is not pointwise bounded on $Z$. To this end, suppose $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$. If $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{e}_{i})|=\infty$ for some $i\in\omega$, then we are done, because $\mathbf{e}_{i}\in Z$. If not, then by the previous lemma, $\left\{f\in\mathcal{I}\colon\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}=\infty\right\}$ is a co-meager subset of $\mathcal{I}$. But $X$ meets every co-meager subset of $\mathcal{I}$, so there is some $f\in X$ with $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f})\big{|}=\infty$. In other words, $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is not pointwise bounded on $X$ (or $Z$). ∎
Theorem 2.3.
There is a barrelled subspace of $\ell^{1}$ with dimension $\leq\!\mathrm{non}(\mathcal{M})$. In particular, $\mathfrak{y}\leq\mathrm{non}(\mathcal{M})$, and it is consistent with $\mathsf{ZFC}$ that $\mathfrak{y}<\mathfrak{c}$.
Proof.
Because $\mathcal{I}$ is homeomorphic to the Baire space $\omega^{\omega}$, which in turn is homeomorphic to $\mathbb{R}\setminus\mathbb{Q}$, there is a non-meager subset of $\mathcal{I}$ with size $\mathrm{non}(\mathcal{M})$. If $X$ is such a set, then $Z=\mathrm{Span}\big{(}\!\left\{\mathbf{e}_{i}\colon i\in\omega\right\}\cup\big{\{}\mathbf{x}_{f}:\,f\in X\big{\}}\big{)}$ is barrelled, by the previous theorem, and $\mathrm{dim}(Z)\leq|X\cup\left\{\mathbf{e}_{i}\colon i\in\omega\right\}|=\mathrm{non}(\mathcal{M})$.
From this, and the definition of $\mathfrak{y}$, it follows immediately that $\mathfrak{y}\leq\mathrm{non}(\mathcal{M})$. The last assertion of the theorem then follows from the fact (well-known to set theorists) that the inequality $\mathrm{non}(\mathcal{M})<\mathfrak{c}$ is consistent with $\mathsf{ZFC}$ (see, e.g., [1]). ∎
We end this section with an observation about rearrangements of the series $\sum_{n\in\omega}\nicefrac{{1}}{{3^{n}}}$. Let us say $\mathbf{x}_{f}\in\ell^{1}$ is a rearrangement of $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$ if $f$ is not merely injective, but bijective.
Observation 2.4.
$\left\{f\in\mathcal{I}\colon f\text{ is a bijection}\right\}$ is a dense $G_{\delta}$ subset of $\mathcal{I}$.
Proof.
For each $n\in\omega$, observe that
$U_{n}\,=\,\left\{f\in\mathcal{I}\colon n\in\mathrm{Image}(f)\right\}\,=\,\textstyle\bigcup\left\{\langle\langle s\rangle\rangle\colon s\text{ is injective and }n\in\mathrm{Image}(s)\right\}$
is a nonempty open subset of $\mathcal{I}$. Given any nonempty open $V\subseteq\mathcal{I}$, fix some basic open $\langle\langle s\rangle\rangle\subseteq V$. If $n\in\mathrm{Image}(s)$ then $\langle\langle s\rangle\rangle\subseteq U_{n}\cap V$. If not, define $t:(|s|+1)\to\omega$ by setting $t\!\restriction\!|s|=s$ and $t(|s|)=n$; then $t$ is injective and $\langle\langle t\rangle\rangle\subseteq U_{n}\cap V$. Either way, $U_{n}\cap V\neq\emptyset$. As $V$ was arbitrary, this shows $U_{n}$ is dense in $\mathcal{I}$. Consequently, $\left\{f\in\mathcal{I}\colon f\text{ is a bijection}\right\}=\bigcap_{n\in\omega}U_{n}$ is a dense $G_{\delta}$ set. ∎
It follows that in the statement of Theorem 2.2, we may, without loss of generality, assume every $f\in X$ is bijective, so that each $\mathbf{x}_{f}$ is a rearrangement. In other words, combining the above observation with the preceding proofs shows:
  • $\circ$
    There is a set $X$ of rearrangements of $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$ such that $|X|=\mathrm{non}(\mathcal{M})$ and $Z=\mathrm{Span}\big{(}\!\left\{\mathbf{e}_{i}\colon i\in\omega\right\}\cup X\big{)}$ is barrelled.
With a little more work, one may also show that if $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{e}_{i})|=\infty$, then $\left\{f\in\mathcal{I}\colon\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{x}_{f})|=\infty\right\}$ is co-meager in $\mathcal{I}$. We omit the details of this, but point out that this observation can be used to improve the above to
  • $\circ$
    There is a set $X$ of rearrangements of $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$ such that $|X|=\mathrm{non}(\mathcal{M})$ and $Z=\mathrm{Span}(X)$ is barrelled.
In other words, Theorem 2.3 is still true for the span of some $\mathrm{non}(\mathcal{M})$ rearrangements of $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$.
Ruckle proved in [14] that for any $\mathbf{x}\in\ell^{1}$, the span of all rearrangements of $\mathbf{x}$ is a barrelled space. This result was also derived as a consequence of a more general theorem by the second author in [20]. The above observation is in this vein, though it neither directly implies the Ruckle result (because it is unclear whether our theorem holds for arbitrary $\mathbf{x}\in\ell^{1}$ in the place of $\left\langle\frac{1}{3^{n}}\colon n\in\omega\right\rangle$), nor is implied by it (because we need only $\mathrm{non}(\mathcal{M})$ rearrangements to get a barrelled span).
3. A generalization
Recall that the density character of a Banach space $X$ is
$d(X)\,=\,\min\left\{|D|\colon D\subseteq X\text{ and }\overline{D}=X\right\},$
i.e., it is the least size of a topologically dense subset of $X$. The topological literature typically refers to this as the density of $X$ (for any space $X$, not just Banach spaces), but the term “density character” is more common in the literature on Banach spaces.
As stated in the introduction, the goal of this section is to generalize the results of Section 2 from $\ell^{1}$ to arbitrary Banach spaces $X$. To do this, we need a set to play the part for $X$ that $\left\{\mathbf{e}_{i}\colon i\in\omega\right\}$ played for $\ell^{1}$. This part will be played by norming sets. Recall that a set of unit vectors $\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}$ in a Banach space $X$ is called norming if for any $\mathbf{y}\in X^{\prime}$, the dual of $X$, $||\mathbf{y}||=\sup_{i\in\kappa}|\mathbf{y}(\mathbf{e}_{i})|$.
Observation 3.1.
An infinite-dimensional Banach space has density character $\kappa$ if and only if it has a norming set of size $\kappa$.
Proof.
On the one hand, if $D\subseteq X$ is a dense subset of $X$ then $\left\{\frac{1}{||\mathbf{d}||}\mathbf{d}\colon\mathbf{d}\in D\setminus\{0\}\right\}$ is a norming set. Thus $X$ has a norming set of size $d(X)$.
On the other hand, suppose $N$ is a norming set for $X$. While a norming set may not be dense in the unit sphere, the span of a norming set is dense in $X$. (This is a consequence of the Hahn-Banach Theorem.) It follows that
$D\,=\,\left\{c_{1}\mathbf{e}_{1}+\dots+c_{n}\mathbf{e}_{k}\colon c_{1},\dots,c_{k}\in\mathbb{Q}\text{ and }\mathbf{e}_{1},\dots,\mathbf{e}_{k}\in N\right\}$
is dense in $X$. And if $N$ is infinite (which must be the case if $X$ is infinite-dimensional), then $|D|=|N|$, and thus $d(X)\leq|N|$. ∎
Following the standard set-theoretic conventions, a cardinal $\kappa$ also denotes a set of size $\kappa$ (the set of all ordinal numbers $<\!\kappa$). Given an infinite cardinal $\kappa$, $\kappa^{\omega}$ denotes the space of functions $\omega\to\kappa$. The topology on $\kappa^{\omega}$ is obtained by giving $\kappa$ the discrete topology and then putting the usual product topology on $\kappa^{\omega}$. This makes $\kappa^{\omega}$ a completely metrizable space of weight $\kappa$, sometimes known as the generalized Baire space of weight $\kappa$ (e.g. in [6]). Just as with the usual Baire space $\omega^{\omega}$, the standard basis for this space consists of sets of the form
$\langle\langle s\rangle\rangle\,=\,\left\{f\in\kappa^{\omega}\colon f\!\restriction\!|s|=s\right\},$
where $s$ is a function $n\to\kappa$ for some $n\in\omega$.
Generalizing the space $\mathcal{I}$ from the previous section, let
$\mathcal{I}_{\kappa}=\left\{f\in\kappa^{\omega}\colon f\text{ is injective}\right\}.$
We claim that $\mathcal{I}_{\kappa}$ is a $G_{\delta}$ subset of $\kappa^{\omega}$. To see this, observe that
$U_{n}\,=\,\left\{x\in\kappa^{\omega}\colon x\!\restriction\!n\text{ is injective}\right\}\,=\,\textstyle\bigcup\left\{\langle\langle s\rangle\rangle\colon|s|=n\text{ and }s\text{ is injective}\right\}$
is open for every $n\in\omega$, and $\mathcal{I}_{\kappa}=\bigcap_{n\in\omega}U_{n}$. By Alexandroff’s Theorem, a subspace of a completely metrizable space is completely metrizable if and only if it is $G_{\delta}$; so in particular, $\mathcal{I}_{\kappa}$ is completely metrizable. (In fact, although this is not really needed here, one can show via [6, Theorem 7.3.8] and [19] that $\mathcal{I}_{\kappa}$ is homeomorphic to $\kappa^{\omega}$ itself. This is analogous to the fact stated in the previous section that $\mathcal{I}$ is homeomorphic to $\omega^{\omega}$.) For convenience, if $s$ is an injective function $n\to\kappa$ then we write $\langle\langle s\rangle\rangle$ rather than $\langle\langle s\rangle\rangle\cap\mathcal{I}_{\kappa}$ to denote the basic open neighborhood of $\mathcal{I}_{\kappa}$
Let $X$ be a Banach space with density character $\leq\!\kappa$, and let $N=\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}$ be (an enumeration of) a norming set in $X$. Given $f\in\mathcal{I}_{\kappa}$, define
$\mathbf{x}_{f}^{N}\,=\,\textstyle\sum_{n\in\omega}\frac{1}{3^{n}}\mathbf{e}_{f(n)}.$
Because $N$ contains only unit vectors, this sum converges and $\mathbf{x}_{f}^{N}\in X$. Thus the map $f\mapsto\mathbf{d}_{f}^{N}$ gives us a natural way to identify $\mathcal{I}_{\kappa}$ with a subset of $X$.
The following lemma is analogous to Lemma 2.1 in the previous section.
Lemma 3.2.
Let $X$ be a Banach space, and let $N=\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}$ be a norming set in $X$. Suppose $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is a sequence in $X^{\prime}$ with $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$, but $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{e}_{i})|<\infty$ for all $i\in\kappa$. Then
$\left\{f\in\mathcal{I}_{\kappa}\colon\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f}^{N})\big{|}=\infty\right\}$
is a co-meager subset of $\mathcal{I}_{\kappa}$.
Proof.
The proof of this lemma is essentially identical to the proof of Lemma 2.1. Only four minor changes need to be made. First, every $\mathcal{I}$ in that proof should be replaced with an $\mathcal{I}_{\kappa}$. Second, every instance of $\mathbf{x}_{f}$ in that proof should be replaced with $\mathbf{x}_{f}^{N}$. Third, in the two paragraphs following the definition of $U_{n}$, we have functions $s:k\to\omega$ and $s_{0}:k\to\omega$; but for the proof of the present lemma these should be functions $k\to\kappa$ instead. Finally, in the paragraph following the definition of $U_{n}$, we should choose $j\in\kappa$ rather than $j\in\omega$, and we should point out that the reason such a choice of $j\in\kappa$ is possible is that $\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}$ is a norming set for $X$. ∎
Theorem 3.3.
Let $X$ be a Banach space, and let $N=\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}$ be a norming set for $X$. If $Y$ is a non-meager subset of $\mathcal{I}_{\kappa}$, then
$Z\,=\,\mathrm{Span}\big{(}\!\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}\cup\big{\{}\mathbf{x}_{f}^{N}:\,f\in Y\big{\}}\big{)}$
is a barrelled subspace of $X$.
Proof.
To show $Z$ is barrelled, it suffices to show that for any sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ in $X^{\prime}$, if $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is pointwise bounded on $Z$ then it is bounded in norm. We show the contrapositive: if a sequence is not bounded in norm, it is not pointwise bounded on $Z$. So suppose $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$. If $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{e}_{i})|=\infty$ for some $i\in\kappa$, then we are done, because $\mathbf{e}_{i}\in Z$. If not, then by the previous lemma, $\left\{f\in\mathcal{I}_{\kappa}\colon\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f}^{N})\big{|}=\infty\right\}$ is a co-meager subset of $\mathcal{I}_{\kappa}$. But $Y$ meets every co-meager subset of $\mathcal{I}_{\kappa}$, so there is some $f\in Y$ such that $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x}_{f}^{N})\big{|}=\infty$. In other words, $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is not pointwise bounded on $X$ (or $Z$). ∎
If $\kappa$ is an infinite cardinal, then $[\kappa]^{\omega}$ denotes the set of all countably infinite subsets of $\kappa$. A family $\mathcal{F}\subseteq[\kappa]^{\omega}$ is cofinal if for every $A\in[\kappa]^{\omega}$ there is some $B\in\mathcal{F}$ such that $B\supseteq A$. In other words, $\mathcal{F}\subseteq[\kappa]^{\omega}$ is cofinal if it is (in the usual sense of the word) cofinal in the poset $([\kappa]^{\omega},\subseteq)$. Define
$\mathrm{cf}[\kappa]^{\omega}\,=\,\min\left\{|\mathcal{F}|\colon\mathcal{F}\text{ is a cofinal subset of }[\kappa]^{\omega}\right\}.$
For example, $\mathrm{cf}[\omega_{1}]^{\omega}=\aleph_{1}$ because, identifying an ordinal with the set of its predecessors as usual, $\left\{\alpha\colon\omega\leq\alpha<\omega_{1}\right\}$ is a cofinal subset of $[\omega_{1}]^{\omega}$. In fact, this example with $\omega_{1}$ can be extended to show that $\mathrm{cf}[\omega_{n}]^{\omega}=\aleph_{n}$ for all $n\in\omega$.
It is not difficult to see that $\kappa\leq\mathrm{cf}[\kappa]^{\omega}$ for any uncountable cardinal $\kappa$. Furthermore, a reasonably straightforward diagonalization argument can be used to show that $\kappa^{+}\leq\mathrm{cf}[\kappa]^{\omega}$ whenever $\kappa$ is a singular cardinal with cofinality $\omega$. Shelah’s Strong Hypothesis, abbreviated $\mathsf{SSH}$, is the statement that these straightforward lower bounds are sharp:
  • $\mathsf{SSH}:$
    $\qquad\qquad\qquad\quad\mathrm{cf}[\kappa]^{\omega}\,=\,\begin{cases}\kappa&\text{ if }\mathrm{cf}(\kappa)>\omega,\\ \kappa^{+}&\text{ if }\mathrm{cf}(\kappa)=\omega.\end{cases}$
Let us note that $\mathsf{SSH}$ is consistent with $\mathsf{ZFC}$. In fact this is true in a rather strong sense: the failure of $\mathsf{SSH}$ implies the consistency of certain large cardinal axioms, and therefore the consistency of $\mathsf{ZFC}$ +¬+\,$\neg$+ $\mathsf{SSH}$ is only provable in strong extensions of $\mathsf{ZFC}$ (see [8] or [3]).
The following analogue of Theorem 2.3 relies on the first author’s recent characterization, in [4], of the size of meager subsets of spaces like $\mathcal{I}_{\kappa}$.
Theorem 3.4.
If $X$ is a Banach space with density character $\kappa$, then $X$ has a barrelled subspace with dimension at most $\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})$.
Proof.
The space $\mathcal{I}_{\kappa}$ is completely metrizable, and it has a basis of size $\kappa$: in fact the standard basis, consisting of the sets of the form $\langle\langle s\rangle\rangle$, has size $\kappa$. By [4, Theorem 3.4], this implies $\mathcal{I}_{\kappa}$ has a non-meager subset $Y$ of size $\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})$. But then $Z=\mathrm{Span}\big{(}\!\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}\cup Y\big{)}$ is barrelled by the previous theorem, and $\mathrm{dim}(Z)\leq\big{|}\!\left\{\mathbf{e}_{i}\colon i\in\kappa\right\}\cup Y\big{|}=\kappa\cdot\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})=\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})$. ∎
Observe that $\mathrm{cf}[\omega]^{\omega}=1$, because the family $\{\omega\}$ is cofinal in $\mathrm{cf}[\omega]^{\omega}$. Thus setting $\kappa=\omega$ in Theorem 3.4 gives an interesting special case:
Corollary 3.5.
Every separable Banach space has a barrelled subspace with dimension $\mathrm{cov}(\mathcal{M})$.
Corollary 3.6.
For every cardinal number $\lambda$, there is a cardinal-preserving notion of forcing which forces $\mathfrak{c}\geq\lambda$ and also forces that every Banach space with density character $<\!\mathfrak{c}$ has a barrelled subspace with dimension $<\!\mathfrak{c}$.
In other words, it is consistent with arbitrarily large values of $\mathfrak{c}$ that for every infinite $\kappa<\mathfrak{c}$, every Banach space with density character $<\!\mathfrak{c}$ has a barrelled subspace with dimension $<\!\mathfrak{c}$.
Proof.
Fix $\lambda$, and fix some $\mu\geq\lambda$ such that $\mathrm{cf}(\mu)>\omega$, and such that $\mathrm{cf}[\kappa]^{\omega}<\mu$ whenever $\kappa<\mu$. (Assuming $\mathsf{SSH}$, this is true for some $\mu\in\{\lambda,\lambda^{+},\lambda^{++}\}$. In $\mathsf{ZFC}$, it is true for any $\beth$-fixed point above $\lambda$ with uncountable cofinality.) Then let $\mathbb{P}$ be the usual notion of forcing to add $\mu$ Cohen reals. Because $\mathbb{P}$ has the ccc, it does not change the value of $\mathrm{cf}[\kappa]^{\omega}$ for any cardinal $\kappa$. Furthermore, $\mathbb{P}$ forces $\mathrm{non}(\mathcal{M})=\aleph_{1}$. By our choice of $\mu$, it follows that $\mathrm{cf}[\kappa]^{\omega}\cdot\mathrm{non}(\mathcal{M})=\mathrm{cf}[\kappa]^{\omega}<\mu$ for any uncountable $\kappa<\mu$ in the forcing extension. Thus the corollary follows from Theorem 3.4. ∎
4. A lower bound for the size of barrelled subspaces
We say that a Banach space $X$ contains a Banach space $Y$ if there is an injective bounded linear map $Y\to X$, or equivalently, if $Y$ is linearly homeomorphic to a (closed) subspace of $X$. In this section we prove that any Banach space whose dual contains either $c_{0}$ or $\ell^{p}$ for some $p\geq 1$ has no infinite-dimensional barrelled subspace with dimension $<\!\mathrm{non}(\mathcal{N})$.
Let $X$ be a Banach space, and let $X^{\prime}$ denote its dual. A sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ in $X^{\prime}$ is norm-unbounded if $\sup_{m\in\omega}||\mathbf{y}_{m}||=\infty$. Let us say that $\mathbf{x}\in X$ detects the norm-unboundedness of a sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ if $\sup_{m\in\omega}\big{|}\mathbf{y}_{m}(\mathbf{x})\big{|}=\infty$. Let us say that $Y\subseteq X$ detects the norm-unboundedness of all functional sequences if for any norm-unbounded sequence $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ in $X^{\prime}$, there is some $\mathbf{x}\in Y$ that detects the norm-unboundedness of $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$.
If $Z\subseteq X$ is barrelled, then $Z$ detects the norm-unboundedness of all functional sequences. (This just restates the property of barrelled spaces given before the proof of Lemma 2.1.) Furthermore, the proof of Theorem 2.2 shows that if $Y\subseteq X$ detects the norm-unboundedness of all functional sequences, then $Z=\mathrm{Span}(Y)$ is barrelled. The following lemma shows the converse is also true.
Lemma 4.1.
Suppose $Z$ is a barrelled subspace of a Banach space $X$. If $Y\subseteq X$ and $Z=\mathrm{Span}(Y)$, then $Y$ detects the norm-unboundedness of all functional sequences.
Proof.
Fix $Y\subseteq X$ with $Z=\mathrm{Span}(Y)$. Our goal is to show that if $Y$ does not detect the norm-unboundedness of every functional sequence, then neither does $Z$.
To this end, suppose $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$ is a norm-unbounded sequence in $X^{\prime}$ such that no $\mathbf{x}\in Y$ detects the norm-unboundedness of $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$. Let $\mathbf{z}\in Z$, and fix $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in Y$ and $c_{1},\dots,c_{n}$ such that $\mathbf{z}=c_{1}\mathbf{x}_{1}+\dots+c_{n}\mathbf{x}_{n}$. For each $i\leq n$, our assumption on $Y$ implies $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{x}_{i})|<\infty$; let $B_{i}=c_{i}\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{x}_{i})|$. Because each of the functionals $\mathbf{y}_{m}$ is linear,
$|\mathbf{y}_{m}(\mathbf{z})|\,=\,\big{|}\sum_{i\leq n}c_{i}\mathbf{y}_{m}(\mathbf{x}_{i})\big{|}\,\leq\,\sum_{i\leq n}c_{i}|\mathbf{y}_{m}(\mathbf{x}_{i})|\,\leq\,\sum_{i\leq n}c_{i}B_{i}.$
Thus $\sup_{m\in\omega}|\mathbf{y}_{m}(\mathbf{z})|\,\leq\,\sum_{i\leq n}c_{i}B_{i}$, which means that $\mathbf{z}$ does not detect the norm-unboundedness of $\left\langle\mathbf{y}_{m}\colon m\in\omega\right\rangle$. ∎
We now move on to proving the main results of this section. The method of proof is different for spaces whose dual contains $\ell^{1}$ and for spaces whose dual contains $c_{0}$ or $\ell^{p}$ for some $p>1$. We handle the $\ell^{1}$ case first.
Lemma 4.2.
Suppose $Y$ is a Banach space containing $\ell^{1}$. Then there is a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of unit vectors in $Y$, and some $\delta>0$, such that
$\big{|}\big{|}\sum_{i=1}^{n}f(i)\mathbf{e}_{i}\big{|}\big{|}\,\geq\,\delta n$
for every function $f:\{1,2,\dots,n\}\to\{-1,1\}$.
Proof.
Let $\phi:\ell^{1}\to Y$ be an injective bounded linear map. By the Open Mapping Theorem, there is some $b>0$ such that $\frac{1}{b}||\mathbf{x}||\leq||\phi(\mathbf{x})||\leq b||\mathbf{x}||$ for all $\mathbf{x}\in\ell^{1}$. For each $i\in\omega$, let $\mathbf{e}^{0}_{i}$ be the sequence in $\ell^{1}$ with a $1$ in coordinate $i$ and a 00 in every other coordinate, and let $\mathbf{e}_{i}=\frac{1}{||\phi(\mathbf{e}_{i}^{0})||}\phi(\mathbf{e}_{i}^{0})$.
Clearly $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ is a sequence of unit vectors in $Y$, and if $c_{1},c_{2},\dots,c_{n}$ is any choice of scalars, then
$\textstyle\big{|}\big{|}\sum_{i=1}^{n}c_{i}\mathbf{e}_{i}\big{|}\big{|}$ $\,=\,\textstyle\big{|}\big{|}\sum_{i=1}^{n}\frac{c_{i}}{||\phi(\mathbf{e}_{i}^{0})||}\phi(\mathbf{e}_{i}^{0})\big{|}\big{|}\,=\,\big{|}\big{|}\phi\big{(}\sum_{i=1}^{n}\frac{c_{i}}{||\phi(\mathbf{e}_{i}^{0})||}\mathbf{e}_{i}^{0}\big{)}\big{|}\big{|}$
$\,\geq\,\textstyle\frac{1}{b}\big{|}\big{|}\sum_{i=1}^{n}\frac{c_{i}}{||\phi(\mathbf{e}_{i}^{0})||}\mathbf{e}_{i}^{0}\big{|}\big{|}\,=\,\textstyle\frac{1}{b}\sum_{i=1}^{n}\frac{|c_{i}|}{||\phi(\mathbf{e}_{i}^{0})||}\,\geq\,\frac{1}{b^{2}}\sum_{i=1}^{n}|c_{i}|.$
In particular, setting $\delta=\frac{1}{b^{2}}$, the lemma just asserts the special case in which $c_{i}=f(i)\in\{-1,1\}$ for all $i\leq n$. ∎
Theorem 4.3.
Let $X$ be a Banach space with the property that its dual space $X^{\prime}$ contains $\ell^{1}$. If $Y\subseteq X$ detects the norm-unboundedness of all functional sequences, then $|Y|\geq\mathrm{cov}(\mathcal{N})$.
Proof.
Suppose $X$ is a Banach space such that $X^{\prime}$ contains $\ell^{1}$. Applying Lemma 4.2, fix a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of unit vectors in $X^{\prime}$ and $\delta>0$ such that for every function $f:\{1,2,\dots,n\}\to\{-1,1\}$, $\big{|}\big{|}\sum_{i=1}^{n}f(i)\mathbf{e}_{i}\big{|}\big{|}\,\geq\,\delta n$.
Our strategy is to define an injective mapping from a probability space $\Omega$ to norm-unbounded sequences in $X^{\prime}$. This gives some meaning to the notion of a “random” norm-unbounded sequence in $X^{\prime}$. We then show that a given $\mathbf{x}\in X$ almost surely does not detect the unboundedness of a “random” sequence in $X^{\prime}$. Once established, this fact can be used to prove the desired inequality.
Let $\Omega=\prod_{i\in\omega}\{-1,1\}$, the space of all functions $\omega\to\{-1,1\}$. We may (and do) view $\Omega$ as a probability space with the standard Lebesgue measure $\mu$, so that
$\mu\big{(}\!\left\{s\in\Omega\colon f(i)=0\right\}\!\big{)}\,=\,\mu\big{(}\!\left\{f\in\Omega\colon s(i)=1\right\}\!\big{)}\,=\,\textstyle\frac{1}{2}$
for all $i\in\omega$. To emphasize the probabilistic viewpoint, let
$P\big{(}\varphi(x)\big{)}\,=\,\mu\big{(}\!\left\{f\in\Omega\colon\varphi(f)\right\}\!\big{)}$
whenever $\varphi(x)$ is some formula/property defining a subset of $\Omega$. The relevant properties of $\mu$ are:
  • $\circ$
    As a probability measure, $\mu$ is countably (sub)additive. In particular,
    $P\big{(}\varphi_{m}(f)\text{ holds for some $m\in A$}\big{)}\,\leq\,\textstyle\sum_{m\in A}P\big{(}\varphi_{m}(f)\big{)}$
    whenever $A\subseteq\mathbb{N}$ and $\varphi_{m}(x)$ is some formula/propety defining a subset of $\Omega$ for every $m\in A$.
  • $\circ$
    The measure space $(\Omega,\mu)$ is an atomless probability Borel measure on a compact metrizable space. All such measure spaces are Borel-isomorphic to the Lebesgue measure on the unit interval $[0,1]$ [9, Theorem 17.41]. In particular, it follows that if $\mathcal{F}$ is a family of $\mu$-null subsets of $\Omega$ and $|\mathcal{F}|<\mathrm{cov}(\mathcal{N})$, then $\bigcup\mathcal{F}\neq\Omega$.
For each $f\in\Omega$ and $m\in\omega$, define
$\mathbf{y}_{m}^{f}\,=\,\textstyle\frac{1}{m^{3}}\sum_{i=1}^{m^{4}}f(i)\mathbf{e}_{i}$
In other words, $\mathbf{y}^{f}_{m}$ is the result of a random walk in $X^{\prime}$, with $m^{4}$ steps, where at step $i$ our walk goes in the direction of either $\mathbf{e}_{i}$ or $-\mathbf{e}_{i}$. For each $f\in\Omega$,
$||\mathbf{y}_{m}^{f}||\,=\,\textstyle\frac{1}{m^{3}}\big{|}\big{|}\sum_{i=1}^{m^{4}}f(i)\mathbf{e}_{i}\big{|}\big{|}\,\geq\,\frac{1}{m^{3}}\delta m^{4}\,=\,\delta m.$
It follows that $\sup_{m\in\omega}||\mathbf{y}^{f}_{m}||=\infty$ for every $f\in\Omega$. In other words, the mapping $f\mapsto\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle$ gives us a way of associating the members of $\Omega$ to norm-unbounded sequences in $X^{\prime}$.
Claim.
For any given $\mathbf{x}\in X$,
$P\big{(}\mathbf{x}\text{ detects the norm-unboundedness of }\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle\!\big{)}\,=\,0.$
Proof of the claim: Fix $\mathbf{x}\in X$ and $m\in\mathbb{N}$. For convenience, let $x_{i}=\mathbf{e}_{i}(\mathbf{x})$ for each $i\leq m^{4}$, so that $\mathbf{y}^{f}_{m}(\mathbf{x})=\frac{1}{m^{3}}\sum_{i=1}^{m^{4}}f(i)x_{i}$ for each $f\in\Omega$. Because $||\mathbf{e}_{i}||=1$, $|x_{i}|\leq||\mathbf{x}||$ for each $i$, and in particular $\sum_{i=1}^{m^{4}}x_{i}^{2}\leq m^{4}||\mathbf{x}||^{2}$.
For each $i\leq m^{4}$, we may (and do) consider the value of each $f(i)x_{i}$ to be a random variable, taking each of the values ±$\pm x_{i}$ with probability $\nicefrac{{1}}{{2}}$. The variance of this random variable is $x_{i}^{2}$. The key fact we need from elementary probability theory is that for independent random variables (which these are), the variance is additive. This means the variance of the value of $\sum_{i=1}^{m^{4}}f(i)x_{i}$ (as $f$ varies) is $\sum_{i=1}^{m^{4}}x_{i}^{2}\leq m^{4}||\mathbf{x}||^{2}$, and therefore standard deviation is $\leq\!m^{2}||\mathbf{x}||$. Furthermore, the mean of $\sum_{i=1}^{m^{4}}f(i)x_{i}$ (as $f$ varies) is 00, by symmetry. By Chebyshev’s inequality,
$P\Big{(}\big{|}\sum_{i=1}^{m^{4}}f(i)x_{i}\big{|}\geq km^{2}||\mathbf{x}||\,\Big{)}\,\leq\,\nicefrac{{1}}{{k^{2}}}$
for any $k\in\mathbb{N}$. Setting $k=m$ and scaling by $\frac{1}{m^{3}}$, it follows that
$P\big{(}|\mathbf{y}^{f}_{m}(\mathbf{x})|\geq{||\mathbf{x}||}\,\big{)}\,=\,P\Big{(}\big{|}\frac{1}{m^{3}}\sum_{i=1}^{m^{4}}f(i)x_{i}\big{|}\geq{||\mathbf{x}||}\,\Big{)}\,\leq\,\nicefrac{{1}}{{m^{2}}}.$
(We note that better bounds are possible, for example via the Azuma-Hoeffding inequality, but the coarser Chebyshev bound suffices to prove the claim.)
Let $\varepsilon>0$. Because $\sum_{m\in\omega}\frac{1}{m^{2}}$ converges, there is some $N\in\mathbb{N}$ such that $\sum_{m\geq N}\frac{1}{m^{2}}<\varepsilon$. Observe that if $m<N$ then
$|\mathbf{y}^{f}_{m}(\mathbf{x})|\,\leq\,||\mathbf{y}^{f}_{m}||||\mathbf{x}||\,\leq\,\delta m||x||\,<\,\delta mN||\mathbf{x}||.$
Combining this with the previous paragraph,
$\textstyle P\big{(}\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|=\infty\big{)}$ $\,\leq\,P\Big{(}\textstyle\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|>\max\big{\{}\delta mN||\mathbf{x}||,{||\mathbf{x}||}\big{\}}\Big{)}$
$\,\leq\,\textstyle P\big{(}|\mathbf{y}_{m}^{f}(\mathbf{x})|>{||\mathbf{x}||}\text{ for some }m\geq N\big{)}$
$\,\leq\,\textstyle\sum_{m\geq N}P\big{(}|\mathbf{y}_{m}^{f}(\mathbf{x})|\geq{||\mathbf{x}||}\big{)}$
$\,<\,\textstyle\sum_{m\geq N}\frac{1}{m^{2}}$
$\,<\,\varepsilon.$
As $\varepsilon$ was arbitrary, this implies $P\big{(}\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|=\infty\big{)}=0$. $\dashv$
To finish the proof of the theorem, let $Y\subseteq X$. Observe that
$\left\{f\in\Omega\colon\text{$Y$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\}$
$\qquad\,=\,\textstyle\bigcup_{\mathbf{x}\in Y}\left\{f\in\Omega\colon\text{$\mathbf{x}$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\}.$
By the claim, this is a size-$|Y|$ union of $\mu$-null sets. I $|Y|<\mathrm{cov}(\mathcal{N})$ then, because $\mu$ is Borel isomorphic to the standard Lebesgue measure on $[0,1]$,
$\bigcup_{\mathbf{x}\in Y}\left\{f\in\Omega\colon\text{$\mathbf{x}$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\}\,\neq\,\Omega.$
In particular, if $|Y|<\mathrm{cov}(\mathcal{N})$ then $Y$ does not detect the norm-unboundedness of every sequence of the form $\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle$, and so $Y$ does not detect the norm-unboundedness of all functional sequences. ∎
Corollary 4.4.
If $X$ is a Banach space whose dual contains $\ell^{1}$, then every barrelled subspace of $X$ has dimension at least $\mathrm{cov}(\mathcal{N})$.
Proof.
Suppose $Z$ is a barrelled subspace of $X$, and fix $Y\subseteq Z$ with $|Y|=\mathrm{dim}(Z)$ and $Z=\mathrm{Span}(Y)$. By Lemma 4.1, $Y$ detects the norm-unboundedness of all functional sequences. By Theorem 4.3, this implies $\mathrm{dim}(Z)=|Y|\geq\mathrm{cov}(\mathcal{N})$. ∎
Corollary 4.5.
It is consistent with $\mathsf{ZFC}$ that no Banach space whose dual contains $\ell^{1}$ has a barrelled subspace with dimension $\mathfrak{b}$.
Proof.
This follows from the previous corollary, together with the fact (well-known to set theorists) that the inequality $\mathfrak{b}<\mathrm{cov}(\mathcal{N})$ is consistent with $\mathsf{ZFC}$. For example, this inequality holds in the so-called “random real model” (see [1]). ∎
Some well known Banach spaces $X$ with the property that $X^{\prime}$ contains $\ell^{1}$ are:
  • $\circ$
    $\ell^{\infty}$, or more generally, $\ell^{\infty}(\kappa)$ for any infinite cardinal $\kappa$,
  • $\circ$
    $c_{0}$ and $c$, or more generally, $c_{0}(\kappa)$ and $c(\kappa)$ for any infinite cardinal $\kappa$,
  • $\circ$
    $L^{\infty}(\Omega,\Sigma,\mu)$ for any non-atomic measure space $(\Omega,\Sigma,\mu)$,
  • $\circ$
    the function space $C(K)$ for any infinite compact Hausdorff space $K$.
Next we prove a result similar to Theorem 4.3, but for Banach spaces having the following general property. Let us say a Banach space $X$ has property ()($\dagger$)( † ) if:
  • ()($\dagger$)( † )
    There is a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of unit functionals in $X^{\prime}$, and a function $f:\omega\to\omega$, such that for every $\mathbf{x}\in X$ with $||\mathbf{x}||=1$,
    |$\big{|}\!\left\{i\in\omega\colon|\mathbf{e}_{i}(\mathbf{x})|>\nicefrac{{1}}{{m}}\right\}\!\big{|}\,<\,f(m)$
    for all sufficiently large values of $m$.
Before going on to how this property relates to barrelledness, let us first show that if $X^{\prime}$ contains $c_{0}$ or $\ell^{p}$ for any $p>1$, then $X$ has property ()($\dagger$)( † ).
Proposition 4.6.
If $X$ is a Banach space whose dual space contains either $c_{0}$ or $\ell^{p}$ for some $p>1$, then $X$ has property ()($\dagger$)( † ).
Proof.
Let us consider first the case where $X^{\prime}$ contains $\ell^{p}$. Fix $p>1$, let $\phi:\ell^{p}\to X^{\prime}$ be an injective bounded linear map, and fix $b>0$ such that $\frac{1}{b}||\mathbf{x}||\leq||\phi(\mathbf{x})||\leq b||\mathbf{x}||$ for all $\mathbf{x}\in\ell^{p}$. For each $i\in\omega$, let $\mathbf{e}_{i}^{0}$ be the sequence in $\ell^{p}$ with a $1$ in coordinate $i$ and a 00 in every other coordinate, and let $\mathbf{e}_{i}=\frac{1}{||\phi(\mathbf{e}^{0}_{i})||}\phi(\mathbf{e}^{0}_{i})$.
Clearly $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ is a sequence of unit vectors in $X^{\prime}$. Fix $\mathbf{x}\in X$ with $||\mathbf{x}||=1$, and define $\varsigma:\mathbb{N}\to\{-1,1\}$ by setting $\varsigma(i)=1$ if $\mathbf{e}_{i}(\mathbf{x})\geq 0$ and $\varsigma(i)=-1$ if $\mathbf{e}_{i}(\mathbf{x})<0$ (i.e., $\varsigma(i)$ indicates the sign of $\mathbf{e}_{i}(\mathbf{x})$). Given $m\in\omega$, let
$F_{m}\,=\,\left\{i\in\omega\colon|\mathbf{e}_{i}(\mathbf{x})|>\nicefrac{{1}}{{m}}\right\},$
and observe that
$\textstyle\frac{1}{m}|F_{m}|$ $\,<\,\textstyle\sum_{i\in F_{m}}|\mathbf{e}_{i}(\mathbf{x})|\,=\,\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}(\mathbf{x})\,\leq\,\big{|}\big{|}\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}\big{|}\big{|}$
$\,\leq\,\textstyle\big{|}\big{|}\sum_{i\in F_{m}}\frac{1}{||\phi(\mathbf{e}_{i}^{0})||}\varsigma(i)\phi(\mathbf{e}_{i}^{0})\big{|}\big{|}\,\leq\,b\big{|}\big{|}\sum_{i\in F_{m}}\varsigma(i)\phi(\mathbf{e}_{i}^{0})\big{|}\big{|}$
$\,\leq\,\textstyle b\big{|}\big{|}\phi\big{(}\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}^{0}\big{)}\big{|}\big{|}\,\leq\,\textstyle b^{2}\big{|}\big{|}\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}^{0}\big{|}\big{|}\,=\,b^{2}|F_{m}|^{\nicefrac{{1}}{{p}}}.$
Consequently, $|F_{m}|^{1-\frac{1}{p}}<mb^{2}$ or, setting $\frac{1}{q}=1-\frac{1}{p}$ as usual, $|F_{m}|<m^{q}b^{2q}$. In particular, taking $f(m)$ to be any integer $>\!m^{q}b^{2q}$ for every $m$, $f$ witnesses ()($\dagger$)( † ).
The argument for $c_{0}$ is almost exactly the same, but with one simplification. In the final equality in the displayed formulas in the second paragraph, we have $b^{2}\big{|}\big{|}\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}^{0}\big{|}\big{|}=b^{2}$ rather than $b^{2}\big{|}\big{|}\sum_{i\in F_{m}}\varsigma(i)\mathbf{e}_{i}^{0}\big{|}\big{|}=b^{2}|F_{m}|^{\nicefrac{{1}}{{p}}}$. Consequently, we get $|F_{m}|<mb^{2}$, and thus taking $f(m)$ to be any integer $>\!mb^{2}$ for every $m$ gives a witness to ()($\dagger$)( † ) in this case. ∎
Lemma 4.7.
Suppose $X$ is a Banach space with property ()($\dagger$)( † ). Then there is a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of unit functionals in $X^{\prime}$, and there is a partition of $\omega$ into finite intervals $I_{1},I_{2},I_{3},\dots$, each of size at least $2$, such that
$\sum_{m=1}^{\infty}\frac{\big{|}\!\left\{i\in I_{m}\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{1}}{{m}}\right\}\!\big{|}}{|I_{m}|}\,<\,\infty$
for every $\mathbf{x}\in X$ with $||\mathbf{x}||=1$.
Proof.
Let $X$ be a Banach space with property ()($\dagger$)( † ), and fix a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of unit functionals in $X^{\prime}$, and a function $f:\omega\to\omega$, witnessing ()($\dagger$)( † ) for $X$. Observe that if $g$ is any function $\omega\to\omega$ strictly larger than $f$, then $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ and $g$ also witness ()($\dagger$)( † ) for $X$ (i.e., increasing $f$ does no harm). Thus, without loss of generality, we may (and do) assume that $f$ is non-decreasing, and that $f(m)\geq 2$ for all $m$.
Let $I_{1},I_{2},I_{3},\dots$ be a partition of $\mathbb{N}$ into adjacent intervals (i.e., $\min I_{1}=1$ and $\max I_{m}=\min I_{m+1}-1$ for all $m$) such that $|I_{m}|>m^{2}f(m)$ for every $m$. By ()($\dagger$)( † ), if $\mathbf{x}\in X$ and $||\mathbf{x}||=1$ then
|$|\left\{i\in I_{m}\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{1}}{{m}}\right\}|\,\leq\,|\left\{i\in\omega\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{1}}{{m}}\right\}|\,<\,f(m)$
for all sufficiently large values of $m$. Consequently, |$\frac{|\left\{i\in I_{m}\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{1}}{{m}}\right\}|}{|I_{m}|}<\nicefrac{{1}}{{m^{2}}}$ for all sufficiently large values of $m$. Because $\sum_{m=1}^{\infty}\nicefrac{{1}}{{m^{2}}}<\infty$, this implies that $\sum_{m=1}^{\infty}\frac{|\left\{i\in I_{m}\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{1}}{{m}}\right\}|}{|I_{m}|}<\infty$ as well. ∎
Theorem 4.8.
Suppose $X$ is a Banach space with property ()($\dagger$)( † ). If $Y\subseteq X$ detects the norm-unboundedness of all functional sequences, then $|Y|\geq\mathrm{cov}(\mathcal{N})$.
Proof.
Suppose $X$ is a Banach space with property ()($\dagger$)( † ). Fix a sequence $\left\langle\mathbf{e}_{i}\colon i\in\omega\right\rangle$ of functionals in $X^{\prime}$, and a sequence $I_{1},I_{2},I_{3},\dots$ of intervals partitioning $\omega$, that witness the conclusion of Lemma 4.7 for $X$.
The basic idea of the proof is the same as for Theorem 4.3: we must define a probability space $\Omega$, and an injective mapping from $\Omega$ to norm-unbounded sequences in $X^{\prime}$, in order to make sense of the notion of a “random” norm-unbounded sequence.
Let $\Omega=\prod_{m\in\omega}I_{m}$, the space of all functions $f:\omega\to\omega$ such that $f(m)\in I_{m}$ for all $m$. In other words, a point of $\Omega$ selects a single member of each $I_{m}$. To make $\Omega$ a probability space, simply consider these selections to happen “randomly” with respect to the uniform distribution on each $I_{m}$. More formally, for each $m$ let $\mu_{m}$ denote the uniform probability measure on $I_{m}$, and then $\mu=\bigotimes_{m\in\omega}\mu_{m}$ is the desired probability measure on $\Omega$. To emphasize the probabilistic viewpoint, let
$P\big{(}\varphi(x)\big{)}\,=\,\mu\big{(}\!\left\{f\in\Omega\colon\varphi(f)\right\}\!\big{)}$
whenever $\varphi(x)$ is some formula/property defining a subset of $\Omega$. The relevant properties of $\mu$ are:
  • $\circ$
    If $F$ is a finite subset of $I_{m}$, then
    $P\big{(}f(m)\in F\big{)}\,=\,\textstyle\nicefrac{{|F|}}{{|I_{m}|}}.$
    This follows from our definition of $\mu$, which uses the uniform probability distribution $\mu_{m}$ on $I_{m}$.
  • $\circ$
    As a probability measure, $\mu$ is countably (sub)additive. In particular,
    $P\big{(}\varphi_{m}(f)\text{ holds for some $m\in A$}\big{)}\,\leq\,\textstyle\sum_{m\in A}P\big{(}\varphi_{m}(f)\big{)}$
    whenever $A\subseteq\mathbb{N}$ and $\varphi_{m}(x)$ is some formula/propety defining a subset of $\Omega$ for every $m\in A$.
  • $\circ$
    The measure space $(\Omega,\mu)$ is an atomless probability Borel measure on a compact metrizable space. (That the measure $\mu$ is atomless follows from our assumption in Lemma 4.7 that $|I_{m}|\geq 2$ for all $m$.) All such measure spaces are Borel-isomorphic to the Lebesgue measure on the unit interval $[0,1]$ [9, Theorem 17.41]. In particular, it follows that if $\mathcal{F}$ is a family of $\mu$-null subsets of $\Omega$ and $|\mathcal{F}|<\mathrm{cov}(\mathcal{N})$, then $\bigcup\mathcal{F}\not=\Omega$.
To each $f\in\Omega$ associate a sequence $\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle$ in $B^{\prime}$ as follows:
$\mathbf{y}^{f}_{m}\,=\,m\mathbf{e}_{f(m)}.$
Observe that $||\mathbf{y}_{m}^{f}||=m||\mathbf{e}_{f(m)}||=m$ for every $f\in\Omega$, hence $\sup_{m\in\omega}||\mathbf{y}^{f}_{m}||=\infty$. Thus, as promised, the map $f\mapsto\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle$ gives us a way of associating the members of our probability space $\Omega$ to norm-unbounded sequences in $X^{\prime}$.
Claim.
For any given $\mathbf{x}\in\ell^{1}$,
$P\big{(}\mathbf{x}\text{ detects the norm-unboundedness of }\left\langle\mathbf{y}^{f}_{m}\colon m\in\omega\right\rangle\!\big{)}\,=\,0.$
Proof of the claim: Fix $\mathbf{x}\in X$, and for convenience, for each $m\in\omega$ let us define $F_{m}=\left\{i\in I_{m}\colon|\mathbf{e}_{i}(\mathbf{x})|\geq\nicefrac{{||\mathbf{x}||}}{{m}}\right\}$. By the first property of $\mu$ listed above,
$P\big{(}|\mathbf{y}^{f}_{m}(\mathbf{x})|\geq||\mathbf{x}||\big{)}\,=\,P\big{(}|\mathbf{e}_{f(m)}(\mathbf{x})|\geq\nicefrac{{||\mathbf{x}||}}{{m}}\big{)}\,=\,\nicefrac{{|F_{m}|}}{{|I_{m}|}}$
for all $m\in\omega$.
[…]
Let $\varepsilon>0$. Because $\sum_{m\in\omega}\nicefrac{{|F_{m}|}}{{|I_{m}|}}$ converges, there is some $N\in\mathbb{N}$ such that $\sum_{m\geq N}\nicefrac{{|F_{m}|}}{{|I_{m}|}}<\varepsilon$. Observe that $|\mathbf{y}^{f}_{m}(\mathbf{x})|\leq||\mathbf{y}^{f}_{m}||||\mathbf{x}||=m||\mathbf{x}||$ whenever $m<N$. Combining this with the observations above and the countable (sub)additivity of our probability measure $\mu$,
$\textstyle P\big{(}\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|=\infty\big{)}$ $\,\leq\,P\big{(}\textstyle\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|>N||x||\big{)}$
$\,\leq\,\textstyle P\big{(}|\mathbf{y}_{m}^{f}(\mathbf{x})|>||\mathbf{x}||\text{ for some }m\geq N\big{)}$
$\,\leq\,\textstyle\sum_{m\geq N}P\big{(}|\mathbf{y}_{m}^{f}(\mathbf{x})|\geq||\mathbf{x}||\big{)}$
$\,=\,\textstyle\sum_{m\geq N}\nicefrac{{|F_{m}|}}{{|I_{m}|}}$
$\,<\,\varepsilon.$
As $\varepsilon$ was arbitrary, this implies $P\big{(}\sup_{m\in\omega}|\mathbf{y}_{m}^{f}(\mathbf{x})|=\infty\big{)}=0$. $\dashv$
To finish the proof of the theorem, let $Y\subseteq X$. Observe that
$\left\{f\in\Omega\colon\text{$Y$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\}$
$\qquad\,=\,\textstyle\bigcup_{\mathbf{x}\in Y}\left\{f\in\Omega\colon\text{$\mathbf{x}$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\},$
and by the claim above, this is a size-$|Y|$ union of $\mu$-null sets. If $|Y|<\mathrm{cov}(\mathcal{N})$, then by the third property of $\mu$ listed above,
$\bigcup_{\mathbf{x}\in Y}\left\{f\in\Omega\colon\text{$\mathbf{x}$ detects the norm-unboundedness of }\left\langle y^{f}_{m}\colon m\in\omega\right\rangle\right\}\,\neq\,\Omega.$
In particular, if $|Y|<\mathrm{cov}(\mathcal{N})$ then $Y$ does not detect the norm-unboundedness of every sequence of the form $\left\langle y^{f}_{m}\colon m\in\omega\right\rangle$, and so $Y$ does not detect the norm-unboundedness of all functional sequences. ∎
Corollary 4.9.
If $X$ is a Banach space with property ()($\dagger$)( † ), then every barrelled subspace of $X$ has dimension at least $\mathrm{cov}(\mathcal{N})$. Moreover, it is consistent with $\mathsf{ZFC}$ that no Banach space with property ()($\dagger$)( † ) has a barrelled subspace of dimension $\mathfrak{b}$.
Proof.
This is proved in exactly the same way as Corollaries 4.4 and 4.5, but using Theorem 4.8 in the place of Theorem 4.3. ∎
Corollary 4.10.
Suppose $X$ is a Banach space whose dual contains either $c_{0}$ or $\ell^{p}$ for some $p\geq 1$. Then every barrelled subspace of $X$ has dimension at least $\mathrm{cov}(\mathcal{N})$. Moreover, it is consistent with $\mathsf{ZFC}$ that no such Banach space has a barrelled subspace of dimension $\mathfrak{b}$.
Proof.
This follows immediately from Proposition 4.6 together with Corollaries 4.4, 4.5, and 4.9. ∎
We end with two questions raised by the results in this section. First, observe that our results apply to an ostensibly larger class of Banach spaces than just those whose dual does not contain $c_{0}$ or $\ell^{p}$ for any $p\geq 1$. Corollary 4.10 really holds for all Banach spaces satisfying the conclusion of Lemma 4.2 or having property ()($\dagger$)( † ). Using ideas from [13], it seems clear that $X^{\prime}$ contains $\ell^{1}$ if and only if $X$ satisfies the conclusion of Lemma 4.2. But it is not clear that property ()($\dagger$)( † ) is equivalent to $X^{\prime}$ containing $c_{0}$ or $\ell^{p}$ for some $p>1$.
Question 4.11.
Does every Banach space either have property ()($\dagger$)( † ) or else have $\ell^{1}$ contained in its dual?
A positive answer to this question would mean that the conclusions of Corollary 4.10 apply to all Banach spaces, which would imply $\mathrm{cov}(\mathcal{N})\leq\mathfrak{y}$, and the consistency of $\mathfrak{b}<\mathfrak{y}$. However, while we suspect the answer to Question 4.11 is negative, we also imagine there could be other roads to proving $\mathrm{cov}(\mathcal{N})\leq\mathfrak{y}$.
Question 4.12.
Is $\mathrm{cov}(\mathcal{N})\leq\mathfrak{y}$?
Question 4.13.
Is $\mathfrak{y}$ equal to any previously studied cardinal characteristic of the continuum, such as $\mathrm{cov}(\mathcal{N})$, $\mathrm{non}(\mathcal{M})$, or $\mathfrak{b}$?